OpenStudy (anonymous):

Abag contains 6 purple marbles, 7 azure marbles and 10 orange marbles. What is the chance of drawing an orange marble? If an orange marble is drawn then placed back into the bag, and a second marble is drawn, what is the probability of drawing an azure marble? Give solutions exactly in reduced fraction form, separated by a comma.

myininaya (myininaya):

there are (6+7+10) marbles in all chance of drawing an orange is 10/(6+7+10)

OpenStudy (anonymous):

total=6+7+10=23 there are 10 orange marbles so chance of drawing an orange marble= 7/23 there 7 azure marbles so, 7/23

myininaya (myininaya):

chance of drawing azure is 7/(6+7+10)